OpenStudy (eric_d):

Problem analysis(equation of motion) A ball is thrown vertically upward with initial velocity of 30 ms^-1. Find a)time taken to reach the highest point b)distance travel

OpenStudy (eric_d):

@Abhisar

OpenStudy (abhisar):

ok, let's see. At the highest point ball will have a zero velocity i.e V=0. U is given 30 m/s\(^2\)

OpenStudy (abhisar):

use the first formula V=U-at and find the time taken, take a=-g since motion is in opp direction of g

OpenStudy (eric_d):

It shld be v=u+at rite

OpenStudy (abhisar):

oh yes,

OpenStudy (eric_d):

3s

OpenStudy (abhisar):

but after taking a=-g it will become v=u-gt

OpenStudy (anonymous):



OpenStudy (anonymous):

If you assume Positive gravity then you also must assume negative velocity because of the rotation of the axis.

OpenStudy (abhisar):

That's correct @eric_d , now use the second equation to find the distance.

OpenStudy (anonymous):